OpenStudy (anonymous):

evaluate the expression x^-2 if x=7

OpenStudy (anonymous):

thats not the way I'm doing it in class i wish it were tho

OpenStudy (anonymous):

thank you anyway

OpenStudy (anonymous):

iThink its 12...

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

negative exponent means reciprocal x^-2 =1/x^2 so \[\frac{ 1 }{ 7^2 }=\frac{ 1 }{ 49 }\]

OpenStudy (anonymous):

oh wait daym hold up .....7^2=49 then 49-2=47

OpenStudy (anonymous):

so it turns out to be 49?

OpenStudy (anonymous):

1/49

OpenStudy (anonymous):

ok i see thank you do u think you could help me with some more problems or no